OpenStudy (anonymous):

how do you simplify this (√x)/x i know that (√x) is x^1/2, but what do i do next? do i move the exponent to the denominator and make it negative?

OpenStudy (cwrw238):

x^1/2 / x = x^1/2 / x^1 = x(1/2 - 1) = x^-1/2 or 1 / x^1/2

OpenStudy (cwrw238):

* last line should be = x^(1/2 - 1) = x^(-1/2) = 1 / x^(1/2)